# Disjointness Violations in Wikidata

### v1

@@ 1 Introduction
This paper examines the role disjointness plays and can play in helping maintain consistency in Wikidata. It describes the current disjointness situation in Wikidata. It examines current issues with disjointness in Wikidata—finding violations and their sources. It describes some reasons why incorrect information causing disjointness violations have ended up in Wikidata. It finally makes suggestions on how to better improve disjointness in Wikidata.
The work described in this paper was performed as part of a larger effort to find and fix issues in the Wikidata ontology.
@@ 4.2 Culprits
A culprit is a highest superclass that has more than one superclass that are disjoint with each other. All culprits are thus disjointness violations. Subclasses of culprits are also disjointness violations but fixing the disjointness on a culprit has a good likelihood of fixing violations on its subclasses, and also on its instances. Culprits are thus good places to start work on fixing disjointness violations.
@@ 4.3 Mistakes and Fixes
Recording the Effectiveness of Fixes
In any systematic investigation of culprits, it is useful to record how many mistakes were identified and how many culprit violations their fix eliminated by checking after each fix which culprits still are disjointness violations. This check can only be approximate as other changes to Wikidata might also fix disjointness violations. [^6] We have not observed any changes to Wikidata that have fixed any significant number of disjointness violations. If only some violations are eliminated this information can be used to show the effectiveness of the work done so far and maybe help with the rest of the process. If all the culprit violations are eliminated this information can be used to show how the culprits are clustered, and maybe point to better ways of identifying related culprits that may then lead to better ways of identifying mistakes and fixes for future rounds of disjointness violation elimination.
@@ 4.4 Example
We have only started the work on identifying mistakes and have not yet performed any fixes so we have only a very little data on effectiveness. Further, we started with culprits that we strongly suspected would lead to mistakes and fixes that eliminate many culprit disjointness violations. We provide here an examination of the most prominent mistake, one that affects the majority of culprits.
@@ 6 Sources of Issues with Disjointness
The above examinations point out and quantify some of the characteristics of disjointness violations in Wikidata. But they do not directly investigate why incorrect information is present in Wikidata, i.e., why there are errors in the Wikidata ontology.
@@ 7 Suggestions for Improving Disjointness
One possibility is to just go through parts of the Wikidata ontology looking for cases where a class appears to be a disjoint union of other classes and adding a “disjoint union of” statement to the class and then determining whether the disjointness is correct based on examining the new disjointness violations.
It is unclear, however, how effective this process would be if performed by users with limited domain knowledge, particularly if the violations can only be detected using QLever with the current week-long delay.
@@
One problem with disjointness in Wikidata is that disjointness violations are not shown to users when they are editing Wikidata unless they use a special plugin—www.wikidata.org/w/index.php?title=User%3ATomT0m%2Fclassification.js&action=raw&ctype=text%2Fjavascript. Making this useful tool a part of the core user interface would help in showing distjointness violations to users.
@@
Disjointness in Wikidata is only possible as part of a disjoint union construct. This can require the creation of artificial classes to be the union.999the class “award, award nominees, award recipients or award ceremony” (Q26877490) is a prime example of an item created to represent a disjointness. The disjoint union construct also requires the use of qualifiers and an artifical item to be the value of the statement. Further, the disjoint classes have to be specified inside the construct, separate from any other information about them.
@@ Acknowledgments
Appendix A Programs and Queries
@@
The following Python program first queries Wikidata to find the disjoint pairs in the disjoint union of statements and then runs a query for each pair to find culprits for the pair.
#!/usr/bin/python
import sys
import os
import subprocess
import csv
import certifi
from SPARQLWrapper import SPARQLWrapper, JSON
# Set SSL certificate path for requests
os.environ[’SSL_CERT_FILE’] = certifi.where()
# Function to execute a SPARQL query using SPARQLWrapper
def execute_sparql_query(sparql_endpoint, query):
    sparql = SPARQLWrapper(sparql_endpoint)
    sparql.setQuery(query)
    sparql.setReturnFormat(JSON)
    print(f"Executing SPARQL query on {sparql_endpoint}...")  # Debug statement
    try:
        results = sparql.query().convert()
        print("SPARQL query executed successfully.")  # Debug statement
        return results
    except Exception as e:
        print(f"Error executing SPARQL query: {e}")
        return None
# Define the SPARQL query for disjoint classes from Wikidata
query_disjointnesses = """
PREFIX wd: <http://www.wikidata.org/entity/>
PREFIX wdt: <http://www.wikidata.org/prop/direct/>
PREFIX p: <http://www.wikidata.org/prop/>
PREFIX pq: <http://www.wikidata.org/prop/qualifier/>
PREFIX wikibase: <http://wikiba.se/ontology#>
PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
SELECT DISTINCT ?class ?classLabel ?l ?e1 ?e1Label ?e2 ?e2Label WHERE {
  ?class p:P2738 ?l .
  MINUS { ?l wikibase:rank wikibase:DeprecatedRank . }
  ?l pq:P11260 ?e1 .
  ?l pq:P11260 ?e2 .
  FILTER ( ( str(?e1) < str(?e2) ) )
  OPTIONAL { ?class rdfs:label ?classLabel . FILTER ( lang(?classLabel) = ’en’ ) }
  OPTIONAL { ?e1 rdfs:label ?e1Label . FILTER ( lang(?e1Label) = ’en’ ) }
  OPTIONAL { ?e2 rdfs:label ?e2Label . FILTER ( lang(?e2Label) = ’en’ ) }
} ORDER BY ?class ?l
# Execute the first query using Wikidata
try:
    print("Starting disjointnesses query...")
    disjointnesses = execute_sparql_query("https://query.wikidata.org/sparql", query_disjointnesses)
    if disjointnesses is None:
        raise Exception("Disjointnesses query returned no results.")
    print("Disjointnesses query completed successfully.")
except Exception as e:
    print(f"Error querying Wikidata: {e}")
    sys.exit(1)
# Initialize the CSV file
print("Initializing CSV files...")
with open("AllCulprits.csv", mode=’w’, newline=’’, encoding=’utf-8’) as file:
    writer = csv.writer(file)
    writer.writerow([
        "Violated Class 1 QID", "Violated Class 1 Label",
        "Violated Class 2 QID", "Violated Class 2 Label",
        "Union Class QID", "Union Class Label",
        "Instance QID", "Instance Label"
with open("Errors_Disjointnesses_Culprits.csv", mode=’w’, newline=’’, encoding=’utf-8’) as file:
    writer = csv.writer(file)
    writer.writerow([
        "Violated Class 1 QID", "Violated Class 1 Label",
        "Violated Class 2 QID", "Violated Class 2 Label",
        "Union Class QID", "Union Class Label"
print("CSV files initialized.")
# Set to track checked combinations
checked_combinations = set()
# QLever API URL
qlever_url = "https://qlever.cs.uni-freiburg.de/api/wikidata"
# Iterate through the results of the first query
print("Processing disjointnesses results...")
for disjointness in disjointnesses["results"]["bindings"]:
    e1 = disjointness.get("e1", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    e1Label = disjointness.get("e1Label", {}).get("value", "")
    e2 = disjointness.get("e2", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    e2Label = disjointness.get("e2Label", {}).get("value", "")
    classEntity = \
          disjointness.get("class", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    classLabel = disjointness.get("classLabel", {}).get("value", "")
    # Print details of the current instance
    print(f"Current instance details - e1: {e1}, e1Label: {e1Label}, e2: {e2}, e2Label: {e2Label},"
          " classEntity: {classEntity}, classLabel: {classLabel}")
    # Skip if this combination has already been checked
    combination = (e1, e2, classEntity)
    if combination in checked_combinations:
        print(f"Skipping already checked combination: {combination}")
        continue
    checked_combinations.add(combination)
    print(f"Processing combination: {combination}")
    # Find culprit instances
    try:
        # Define the SPARQL query for finding culprits using QLever
        culprits_query = f"""
        PREFIX violated_1: <http://www.wikidata.org/entity/{e1}>
        PREFIX violated_2: <http://www.wikidata.org/entity/{e2}>
        PREFIX wdt: <http://www.wikidata.org/prop/direct/>
        PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
        SELECT DISTINCT ?class ?classLabel
        WHERE {{
            SELECT ?class
            WHERE {{
              ?class wdt:P279+ violated_1: .
              ?class wdt:P279+ violated_2: .
          MINUS {{
            ?class wdt:P279 ?parent .
            ?parent wdt:P279+ violated_1: .
            ?parent wdt:P279+ violated_2: .
          OPTIONAL {{ ?class rdfs:label ?classLabel . FILTER ( lang(?classLabel) = ’en’ ) }}
        # Set up curl command for QLever
        print("Preparing curl command for QLever...")
        commandQ = [
            ’curl’, ’-s’, qlever_url,
            ’-H’, ’Accept: text/tab-separated-values’,
            ’-H’, ’Content-type: application/sparql-query’,
            ’--data’, culprits_query
        print(f"Executing query with curl: {culprits_query[:80]}...")  # Log the start of the query
        try:
            culprits_result = subprocess.check_output(commandQ, text=True)
            print("QLever query execution successful.")  # Debug statement
        except subprocess.CalledProcessError as e:
            print(f"Error executing QLever query: {e}")
            culprits_result = None
        if culprits_result is None:
            print(f"No culprits found for combination: {combination}")
            continue
        # Process the culprits results
[…]
        if len(culprits_lines) > 1:
            for line in culprits_lines[1:]:
[…]
                instance = columns[0].replace("http://www.wikidata.org/entity/", "")
                instanceLabel = columns[1] if len(columns) > 1 else ""
                print(f"Found instance: {instance}, Label: {instanceLabel}")
                # Output to CSV
                with open("AllCulprits.csv", mode=’a’, newline=’’, encoding=’utf-8’) as file:
                    writer = csv.writer(file)
                    writer.writerow([e1, e1Label, e2, e2Label, classEntity, classLabel,
                                         instance, instanceLabel])
                print(f"Logged instance to CSV: {instance}")
    except Exception as e:
        print(f"Error processing combination {combination}: {e}")
        # Log the error to Errors_Disjointnesses_Culprits.csv
        with open("Errors_Disjointnesses_Culprits.csv", mode=’a’, newline=’’, encoding=’utf-8’) as file:
            writer = csv.writer(file)
            writer.writerow([e1, e1Label, e2, e2Label, classEntity, classLabel])
print("Processing completed.")
The following Python program first queries Wikidata to find the disjoint pairs in the disjoint union of statements and then runs a query for each pair that finds the number of culprits for the pair and the number of disjointness violations at both the subclass and instance level. It produces one line of output for each disjointness pair with violations.
#!/bin/python
import os
import subprocess
import csv
import certifi
from SPARQLWrapper import SPARQLWrapper, JSON
# Set SSL certificate path for requests
os.environ[’SSL_CERT_FILE’] = certifi.where()
# Function to execute a SPARQL query using SPARQLWrapper
def execute_sparql_query(sparql_endpoint, query):
    sparql = SPARQLWrapper(sparql_endpoint)
    sparql.setQuery(query)
    sparql.setReturnFormat(JSON)
    print(f"Executing SPARQL query on {sparql_endpoint}...")  # Debug statement
    try:
        results = sparql.query().convert()
        print("SPARQL query executed successfully.")  # Debug statement
        return results
    except Exception as e:
        print(f"Error executing SPARQL query: {e}")
        return None
# Define the SPARQL query for disjoint classes from Wikidata
query_disjointnesses = """
PREFIX wd: <http://www.wikidata.org/entity/>
PREFIX wdt: <http://www.wikidata.org/prop/direct/>
PREFIX p: <http://www.wikidata.org/prop/>
PREFIX pq: <http://www.wikidata.org/prop/qualifier/>
PREFIX wikibase: <http://wikiba.se/ontology#>
PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
SELECT DISTINCT ?class ?classLabel ?l ?e1 ?e1Label ?e2 ?e2Label WHERE {
  ?class p:P2738 ?l .
MINUS { ?class wdt:P279* wd:Q573 . } # day
  MINUS { ?l wikibase:rank wikibase:DeprecatedRank . }
  ?l pq:P11260 ?e1 .
  ?l pq:P11260 ?e2 .
  FILTER ( ( str(?e1) < str(?e2) ) )
  OPTIONAL { ?class rdfs:label ?classLabel . FILTER ( lang(?classLabel) = ’en’ ) }
  OPTIONAL { ?e1 rdfs:label ?e1Label . FILTER ( lang(?e1Label) = ’en’ ) }
  OPTIONAL { ?e2 rdfs:label ?e2Label . FILTER ( lang(?e2Label) = ’en’ ) }
} ORDER BY ?class ?l
# Execute the first query using Wikidata
try:
    print("Starting disjointnesses query...")
    disjointnesses = execute_sparql_query("https://query.wikidata.org/sparql", query_disjointnesses)
    if disjointnesses is None:
        raise Exception("Disjointnesses query returned no results.")
    print("Disjointnesses query completed successfully.")
except Exception as e:
    print(f"Error querying Wikidata: {e}")
    sys.exit(1)
# Initialize the CSV file
print("Initializing CSV file...")
with open("AllNumbers.csv", mode=’w’, newline=’’, encoding=’utf-8’) as file:
    writer = csv.writer(file)
    writer.writerow([
        "Violated Class 1 QID", "Violated Class 1 Label",
        "Violated Class 2 QID", "Violated Class 2 Label",
        "Union Class QID", "Union Class Label",
        "Number of Culprits", "Number of Classes", "Number of Instances",
        "Number of Classes and Instances"
print("CSV file initialized.")
# Set to track checked combinations
checked_combinations = set()
# Initialize the cycle for 0 results printing
cycle_symbols = ["-", "/", "|", "\\"]
cycle_index = 0
# QLever API URL
qlever_url = "https://qlever.cs.uni-freiburg.de/api/wikidata"
# Iterate through the results of the first query
print(f"Processing {len(disjointnesses["results"]["bindings"])} disjointnesses results...")
for disjointness in disjointnesses["results"]["bindings"]:
    e1 = disjointness.get("e1", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    e1Label = disjointness.get("e1Label", {}).get("value", "")
    e2 = disjointness.get("e2", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    e2Label = disjointness.get("e2Label", {}).get("value", "")
    classEntity = \
        disjointness.get("class", {}).get("value", "").replace("http://www.wikidata.org/entity/", "")
    classLabel = disjointness.get("classLabel", {}).get("value", "")
    # Skip if this combination has already been checked
    combination = (e1, e2, classEntity)
    if combination in checked_combinations:
        print(f"Skipping already checked combination: {combination}")
        continue
    checked_combinations.add(combination)
    print(f"Processing combination: {combination} e1: {e1Label} e2: {e2Label} CLASS: {classLabel}")
    # Find culprit instances
    try:
        # Define the SPARQL query for finding culprits using QLever (without the wikibase prefix)
        culprits_query = f"""
                PREFIX violated_1: <http://www.wikidata.org/entity/{e1}>
                PREFIX violated_2: <http://www.wikidata.org/entity/{e2}>
                PREFIX wdt: <http://www.wikidata.org/prop/direct/>
                SELECT (COUNT(?class) AS ?count)
                WHERE {{
                    SELECT ?class
                    WHERE {{
                      ?class wdt:P279+ violated_1: .
                      ?class wdt:P279+ violated_2: .
                  MINUS {{
                    ?class wdt:P279 ?parent .
                    ?parent wdt:P279+ violated_1: .
                    ?parent wdt:P279+ violated_2: .
        # Set up curl command for QLever
        commandQ = [
            ’/usr/bin/curl’, ’-s’, qlever_url,
            ’-H’, ’Accept: text/tab-separated-values’,
            ’-H’, ’Content-type: application/sparql-query’,
            ’--data’, culprits_query
        try:
            culprits_result = subprocess.check_output(commandQ, text=True)
        except subprocess.CalledProcessError as e:
            print(f"Error executing QLever query: {e}")
            culprits_result = None
        if culprits_result is None:
            print(f"No culprits found for combination: {combination}")
            continue
        # Process the culprits results
[…]
        if len(culprits_lines) > 1:
            culprit_count = culprits_lines[1].strip()
            print(f"Found {culprit_count} culprits for combination: {combination}")
            # Additional Queries for Class and Instance Counts
            # Query to get the number of all classes related to the disjointness
            classes_query = f"""
            PREFIX wd: <http://www.wikidata.org/entity/>
            PREFIX wdt: <http://www.wikidata.org/prop/direct/>
            PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
            SELECT (COUNT(DISTINCT ?class) AS ?classCount) WHERE {{
              {{ ?class wdt:P279* wd:{e1} . }}
              UNION
              {{ ?class wdt:P279* wd:{e2} . }}
            # Execute the classes count query
            print(f"Executing classes count query for combination: {combination}")
            command_classes_count = [
                ’/usr/bin/curl’, ’-s’, qlever_url,
                ’-H’, ’Accept: text/tab-separated-values’,
                ’-H’, ’Content-type: application/sparql-query’,
                ’--data’, classes_query
            try:
                classes_count_result = subprocess.check_output(command_classes_count, text=True)
[…]
                classes_count = classes_count_lines[1].strip() if len(classes_count_lines) > 1 else "0"
                print(f"Found {classes_count} classes for combination: {combination}")
            except subprocess.CalledProcessError as e:
                print(f"Error executing classes count query: {e}")
                classes_count = "0"
            # Query to get the number of all classes and instances related to the disjointness
            instances_query = f"""
            PREFIX wd: <http://www.wikidata.org/entity/>
            PREFIX wdt: <http://www.wikidata.org/prop/direct/>
            PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>
            SELECT (COUNT(DISTINCT ?entity) AS ?entityCount) WHERE {{
              {{ ?entity wdt:P31/wdt:P279* wd:{e1} . }}
              UNION
              {{ ?entity wdt:P31/wdt:P279* wd:{e2} . }}
            # Execute the classes and instances count query
            print(f"Executing classes and instances count query for combination: {combination}")
            command_instances_count = [
                ’/usr/bin/curl’, ’-s’, qlever_url,
                ’-H’, ’Accept: text/tab-separated-values’,
                ’-H’, ’Content-type: application/sparql-query’,
                ’--data’, instances_query
            try:
                instances_count_result = subprocess.check_output(command_instances_count,
                                                                             text=True)
[…]
                instances_count = instances_count_lines[1].strip() if len(
                    instances_count_lines) > 1 else "0"
                print(f"Found {int(classes_count)} {int(instances_count)} {int(classes_count) +"
                      " int(instances_count)} classes and instances for combination: {combination}")
            except subprocess.CalledProcessError as e:
                print(f"Error executing classes and instances count query: {e}")
                instances_count = "0"
            # Output to CSV
            with open("AllNumbers.csv", mode=’a’, newline=’’, encoding=’utf-8’) as file:
                writer = csv.writer(file)
                writer.writerow([
                    e1, e1Label, e2, e2Label, classEntity, classLabel,
                    culprit_count, classes_count, instances_count,
                    int(classes_count) + int(instances_count)
            print(f"Logged to CSV: {culprit_count}, {classes_count}, instances_count, "
                   "{int(classes_count) + int(instances_count)}")
        else:
[…]
            cycle_index = (cycle_index + 1) % 4
    except Exception as e:
        print(f"Error processing combination {combination}: {e}")
print("Processing completed.")

### v2

@@ 1 Introduction
This paper examines the role disjointness plays and can play in helping maintain consistency in Wikidata. It describes the current disjointness situation in Wikidata. It examines current issues with disjointness in Wikidata—finding violations and their sources. It describes some reasons why incorrect information causing disjointness violations have ended up in Wikidata. It finally makes suggestions on how to better improve disjointness in Wikidata. The work described in this paper was performed as part of a larger effort to find and fix issues in the Wikidata ontology.
@@ 4.2 Culprits
If a class $A$ forms a disjointness violation because it is a subclass of both classes in a disjointness pair, then all its subclasses will also be disjointness violations because they are also subclasses of both of the disjoint classes. Fixing the violation for $A$ will likely fix the disjointness violations for its subclasses. It is thus useful to first look at the most-general classes that form disjointness violations for each disjointness pair. We call these most-general classes culprits.
For example, “bow” is a culprit because it is a subclass of both elements of the disjointness pair “gun” and ”draft weapon” and none of its superclasses are subclass of both of the disjoint classes. The violations for the 26 subclasses of “bow” will very likely be fixed if the violation on “bow” is fixed.
@@ 4.3 Mistakes and Fixes
In any systematic investigation of culprits, it is useful to record how many mistakes were identified and how many culprit violations their fix eliminated by checking after each fix which culprits still are disjointness violations. This check can only be approximate as other changes to Wikidata might also fix disjointness violations. [^6] We have not observed any other changes to Wikidata that have fixed any significant number of disjointness violations. If only some violations are eliminated this information can be used to show the effectiveness of the work done so far and maybe help with the rest of the process. If all the culprit violations are eliminated this information can be used to show how the culprits are clustered, and maybe point to better ways of identifying related culprits that may then lead to better ways of identifying mistakes and fixes for future rounds of disjointness violation elimination.
@@ 4.4 Example
We have only started the work on identifying mistakes so we have only a little data on effectiveness. Our identification of mistakes started with culprits that we strongly suspected would lead to mistakes and fixes that eliminate many culprit disjointness violations. We provide here an examination of the most prominent mistake, one that affects the majority of culprits.
@@ 6 Sources of Issues with Disjointness
The above examinations point out and quantify some of the characteristics of disjointness violations as they occur in Wikidata. But they do not directly speak to why this incorrect information is present in Wikidata, i.e., what is the reason that Wikidata users added information to Wikidata that created disjointness violations.
@@ 7 Suggestions for Improving Disjointness
One possibility is to just go through parts of the Wikidata ontology looking for cases where a class appears to be a disjoint union of other classes and adding a “disjoint union of” statement to the class and then determining whether the disjointness is correct based on examining the new disjointness violations. It is unclear, however, how effective this process would be if performed by users with limited domain knowledge, particularly if the violations can only be detected using QLever with the current week-long delay.
@@
One problem with disjointness in Wikidata is that disjointness violations are not shown to users when they are editing Wikidata unless they use a special plugin—www.wikidata.org/w/index.php?title=User%3ATomT0m%2Fclassification.js&action=raw&ctype=text%2Fjavascript. Making this useful tool a part of the core user interface would help in showing disjointness violations to users.
@@
Disjointness in Wikidata is only possible as part of a disjoint union construct. This can require the creation of artificial classes to be the union, such as “award, award nominees, award recipients or award ceremony” (Q26877490). The disjoint union construct also requires the use of qualifiers and an artifical item to be the value of the statement. Further, the disjoint classes have to be specified inside the construct, separate from any other information about them.
@@ 8 Summary and Future Work
We are continuing to fix disjointness violations. The current state of our work is described at www.wikidata.org/wiki/User:Egezort/Fixing_all_(yes_all)_Disjointness_Violations. We have tackled most of the disjointness statements that produced a small number of violations, leaving the larger ones for later, perhaps as a community effort.
@@ Acknowledgments
Appendix A Programs and Results
@@
The Python program AllCulprits.py in the ancillary files first queries Wikidata to find the disjoint pairs in the disjoint union of statements and then runs a query for each pair to find culprits for the pair.
The main results are in AllCulprits.csv, showing the two disjoint classes of the pair, the class that the disjoint union is on, and a culprit class for the pair. Each class is shown as its identifier in Wikidata and its English label, if any. The violations between “concrete object” and “abstract entity” with labels starting with LOC are for genes.
The program also outputs an error file, but no errors were encountered.
The Python program AllNumbers.py in the ancillary files first queries Wikidata to find the disjoint pairs in the disjoint union of statements and then runs a query for each pair that finds the number of culprits for the pair and the number of disjointness violations at both the subclass and instance level.
The main results are in AllNumbers.csv, which contains one line of output for each disjointness pair that has violations, shown as for AllCulprits.py, with the number of culprits for the pair, the total number of subclasses of all the culprits for the pair, the total number of instances of all the culprits for the pair, and the total of both subclasses and instances of all the culprits for the pair.
The program also outputs all disjoint pairs, even those with no violations, in AllDisjoints.csv.